Buttermilk has 6% butterfat, 1% milk has (obviously) 1% butterfat. How much buttermilk must be added to 5 gallons of 1% milk to make a solution of "whole milk" (whole milk has 4% buttermilk)?
Answer by mananth(16946)

------ percent ---------------- quantity
1% milk 1 ----------------5 gallons
Buttermilk 6----------------x gallons
Total 4---------------- 5+x gallons
1*5+6*x=4(5+x)
5+6x=20+4x
6x-4x=20-5
2x=15
/2
x=7.5 gallons Buttermilk